OpenStudy (anonymous):

What is the simplest form of the expression? 3sqrt(750)+3sqrt(2058)-3sqrt(48) a. 14 3sqrt(6) b. 17.5 3sqrt(6) c. 33 3sqrt(6) d. 10 3sqrt(6)

OpenStudy (anonymous):

\[14 \sqrt[3]{6} , 17.5 \sqrt[3]{6}, 33 \sqrt[3]{6}, 10 \sqrt [3]{6}\]

OpenStudy (anonymous):

\[\sqrt[3]{750} + \sqrt[3]{2058} - \sqrt[3]{48}\]

OpenStudy (mathstudent55):

Let's start with the prime factors of 750, 2058, and 48 750/2 = 375 375/3 = 125 125/5 = 25 25/5 = 5 5/5 = 1 750 = 2 x 3 x 5^3 2058/2 = 1029 1029/3 = 343 343/7 = 49 49/7 = 7 7/7 = 1 2058 = 2 x 3 x 7^3 48/2 = 24 24/2 = 12 12/2 = 6 6/2 = 3 3/3 = 1 48 = 2^4 x 3

OpenStudy (anonymous):

ok yes

OpenStudy (mathstudent55):